OpenStudy (anonymous):

@jeremyggg Simplify 2(x + 4) + 3(x - 4) this is my next question lol im not good at simplifing at all can you walk me through it?

OpenStudy (anonymous):

ah ok! it worked!

OpenStudy (igreen):

\(2(x + 4) + 3(x - 4)\) You will distribute 2 into the first parenthesis by multiplying it to 'x' and '4'. \(2 \times x = 2x\) \(2 \times 4 = 8\) Now we have: \('2x + 8'\). Then we distribute 3 into the second parenthesis by multiplying it to 'x' and '-4'. \(3 \times x = 3x\) \(3 \times -4 = -12\) So we have \('3x - 12'\). Now we combine our two equations and keep the addition sign: \((2x + 8) + (3x - 12)\)

OpenStudy (anonymous):

ok now how do i do that last part so i can get my answer?

OpenStudy (igreen):

The last part is combining like terms: This is where we left off: \(2x + 8 + 3x - 12\) Our like terms are: \((2x, 3x)\), and \((8, -12)\) \(2x + 3x = 5x\) \(8 - 12 = -4\) So your final answer is \('5x - 4'\).

OpenStudy (igreen):

Do you get it? @stocksro

OpenStudy (anonymous):

yes thats how i thought i sould do it im trying to do my best on this pre test so i can pass algebra 2

OpenStudy (igreen):

Okay.

OpenStudy (anonymous):

2(x + 4) + 3(x - 4) ahh would you like a little teaching on how to never get confused when simplifying? a little understanding of how to control these signs?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

if there is ever a number multiplied by another number you will do this math RIGHT away. if i said x = 2 and said what is 2x +3? you could solve this easy right?

OpenStudy (anonymous):

now... we are gonna make it a little complex for a second, but just trust me! what about 2(x+1) + 3? at the moment we dont know what x is! x+1 equals some number that we cant know because x isnt given so lets just say x+1 is y. now it is like saying 2y + 3!

OpenStudy (anonymous):

we arent taking this approach to solve this because it is a bad idea to complicate it! i just wanted to explain how 2(x+1)+3 and 2x+3 are EXACTLY alike.

OpenStudy (anonymous):

lol god i get so confused i math

OpenStudy (anonymous):

now... onto the problem! 2(x + 4) + 3(x - 4) what do we do first? we look for what is being multiplied and separated by the plus sign. I like to think of the plus sign as the final bit of work to be done last. if you REALLY wanna not be confused... put brackets around everything that the plus sign separates! (dont worry, this will all come together soon!)

OpenStudy (anonymous):

[2(x+4)] + [3(x+4)]

OpenStudy (anonymous):

now that these are separated by the brackets we can do one side at a time! you didnt get lost here right?

OpenStudy (anonymous):

nope im still with you

OpenStudy (anonymous):

ok so let us not lose these brackets! ill do the left side of the + you do the right once i finish.

OpenStudy (anonymous):

[2(x+4)] i distribute this 2!

OpenStudy (anonymous):



OpenStudy (anonymous):

keeping up with my brackets! nothing can leave them.

OpenStudy (anonymous):

nooo i messed

OpenStudy (anonymous):

[2x+8] is my side

OpenStudy (anonymous):

[2(x+4)] + [3(x+4)] now do your side

OpenStudy (anonymous):

3x+12 right?

OpenStudy (anonymous):

yep but dont lose the brackets until we are done with as much as we can!

OpenStudy (anonymous):

which.. at this point we are.

OpenStudy (anonymous):

[2x+8] + [3x+12] now we are ready to remove em and simplify. (this goes much faster once you understand!) 2x+8 +3x+12 without the brackets. you can take it from here?

OpenStudy (anonymous):

no wait!

OpenStudy (anonymous):

i copied the problem wrong! sorry. its 3x-12 because your side is [3(x-4)] oops! so 2x+8+3x-12

OpenStudy (anonymous):

i hope you have another one of these because i bet you can solve it super fast the second time around

OpenStudy (anonymous):

so the answer is 5x-4?

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

you are!

OpenStudy (anonymous):

my approach when i did these back then, give a little space between those plus signs and distribute right away for the parenthesis. then just do what you did to get 5x-4!

OpenStudy (anonymous):

now this is another question that i need you to walk me through

OpenStudy (anonymous):

im ready!

OpenStudy (anonymous):

If g(x) = 2x + 2, find g(a + h) - g(a).

OpenStudy (anonymous):

and then it gives you a and h right?

OpenStudy (anonymous):

and btw i had another one that was like this Simplify 3(x - 2) + 5x i got 8x-6 did i do it right :) and no that was the whole problem then it gives me 3 answer choices and those are 2h 2h+4 h+2

OpenStudy (anonymous):

Yes that is right!

OpenStudy (anonymous):

If g(x) = 2x + 2, find g(a + h) - g(a).

OpenStudy (anonymous):

ok this one is set up to sound tricky. but it also isn't!

OpenStudy (anonymous):

lol great i understand how to simplify !!! and how do i do this tricky problem

OpenStudy (anonymous):

the minus sign is just an unfancy plus sign! give them some space also. so... g(a+h) - g(a)

OpenStudy (anonymous):

we know that g(x) =2x+2 so this is our formula we use. let's see you explain this one! do the left side first!

OpenStudy (anonymous):

remember, we just plug in for x whatever is in the parenthesis!

OpenStudy (anonymous):

ooh and when you sub for x, everything being subbed must be in parenthesis where the x was.

OpenStudy (anonymous):

g(x) =2(x)+2

OpenStudy (anonymous):

think you can do it?

OpenStudy (anonymous):

g(x) = 2x+2. FIND g(a+h) - g(a)

OpenStudy (anonymous):

can you do this one bc im so lost in it

OpenStudy (anonymous):

ill tell you what to do and you solve it! i think you'll be surprised

OpenStudy (anonymous):

first, lets work on the left side! g(x) = 2x+2 g(a+h) - g(a) you gotta plug in what is in the parenthesis for x!

OpenStudy (anonymous):

here. ill do the right side first. g(a) so if g(x) =2x+2 g(a) = 2(a) +2

OpenStudy (anonymous):

so you just took a and put it in for x

OpenStudy (anonymous):

yep. just like we do with numbers. nothing is different or scary with letters! letters are easier than numbers!

OpenStudy (anonymous):

g(a+h) - g(a) i found g(a) now you find g(a+h)!

OpenStudy (anonymous):

g(x) =2x+2

OpenStudy (anonymous):

so it would be 2(a+h)+2

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

now simplify that!

OpenStudy (anonymous):

can you show me bc the +h throws me off

OpenStudy (anonymous):